OpenStudy (anonymous):

if 3y+3=x-11 and 5y-6=x+2 what do variables x and y equal

OpenStudy (anonymous):

3y + 3 = x - 11 3y - x = -11 - 3 3y -x = - 14 <-- remember this 5y - 6 = x + 2 5y - x = 2 + 6 5y - x = 8 < -- remember this now take the two equations that I told you to remember... 3y - x = - 14--> (-1)3y - x = - 14 5y - x = 8 ------------- -3y + x = 14 (result of multiplying by -1) 5y - x = 8 -------------add 2y = 22 y = 11 now sub 11 in for y in either of the two equations 5y - x = 8 5(11) - x = 8 55 - x = 8 -x = 8 - 55 -x = - 47 x = 47 check... 3y + 3 = x - 11 3(11) + 3 = 47 - 11 33 + 3 = 47 - 11 36 = 36 (correct) check... 5y - 6 = x + 2 5(11) - 6 = 47 + 2 55 - 6 = 49 49 = 49 (correct) ANSWER : x = 47 and y = 11 or (47,11)